OpenStudy (anonymous):

Hi I need to write the equations for the following please help: I also have done the solutions against them, please check them and let me know if they are right: 1. the sum of multiplicative inverses of two positive even integers is 5/12. Find the integers. 2. Find the natural numbers which when added to its square gives 9 times the next natural number.

OpenStudy (anonymous):

3. Joe bought some chairs for Rs. 200. 5 chairs were stolen from the total. He sold the remaining chairs for a profit of Rs. 40. The selling price of each chair was Rs. 4 more than its cost price. How many chairs did he buy and how much was the cost price of each chair?

OpenStudy (anonymous):

4. The denominator of a fraction is 4 more than its numerator. Taking the square of this fraction we get a new fraction. The difference between the denominator and numerator of the new fraction is 40. Find the fraction.

OpenStudy (anonymous):

Here are my answers now: 1. \[(1/x) - (1/x+2) = 5/12\]

OpenStudy (turingtest):

what is Rs ?

OpenStudy (anonymous):

2. x \[x + x ^{2} = 9 (x+1)\]

OpenStudy (turingtest):

1. is wrong 2. is right ...

OpenStudy (anonymous):

Rs. (Rupees) is a currency similar to Dollars.

OpenStudy (turingtest):

Sorry, should have known.

OpenStudy (turingtest):

what is the multiplicitive inverse of a number?

OpenStudy (anonymous):

If x is a number, it's multiplicative inverse is 1/x

OpenStudy (turingtest):

right, so we have two positive even integers. But it doesn't say "successive" so we can't write them as x and x+2 it also says "sum" not "difference" can you fix it do you think?

OpenStudy (anonymous):

Oh yeah I wrote the minus sign by mistake, but I'm unable to figure out how to express those two even integers?

OpenStudy (turingtest):

Start off calling the two numbers 'a' and 'b'. that gives us\[\frac{1}{a}+\frac{1}{b}=\frac{5}{12}\]but we know that these numbers are even, so we can write them as. a=2n b=2m because any even number is 2 times SOME integer. So what will you get after using that?

OpenStudy (anonymous):

1/2n + (1/2m) = 5/12 But then how would we solve the answer for x?

OpenStudy (turingtest):

we don't it would seem do you know that you need to solve it?

OpenStudy (anonymous):

Yes, i have to make the equations, and then solve for x. The answer given at the end of the book for the question is 4 and 6.

OpenStudy (anonymous):

I think 2x and 2x+2 as the even integers, even if they donts say successive even numbers, we dont have any other solution for it.

OpenStudy (anonymous):

dont*

OpenStudy (turingtest):

This is what is called a Diophantine equation. That means the answers can only be integers and the number of variables is more than the number of solutions. Since the answers are 4 and 6 that would suggest that they meant to write the word 'successive' and didn't. Seems to be a mistake in the book.

OpenStudy (turingtest):

If they mean successive then your formula is correct (with the + fixed). Note that if we call them 2x and 2x+2 we need to multiply x by 2 at the end to find the actual numbers.

OpenStudy (anonymous):

Yes, definitely seem sto be an error in the book. Ok, now for the third one, i'm pretty clueless.

OpenStudy (turingtest):

gotta read it one sec...

OpenStudy (turingtest):

Ok first we need a formula for how much he spent. Do you have that?

OpenStudy (anonymous):

He spent $ 200.

OpenStudy (anonymous):

He spent $ 200.

OpenStudy (anonymous):

he bought x chairs, 5 weres tolen so, x-5

OpenStudy (turingtest):

you know, again the wording is weak is it 200 total? or 200 per chair? it's rupees so I thought 200 per chair, right?

OpenStudy (anonymous):

He spent $ 200.

OpenStudy (anonymous):

No, "Joe bought some chairs for $ 200"

OpenStudy (anonymous):

He spent $ 200.

OpenStudy (anonymous):

so "some" is definitely "x"

OpenStudy (turingtest):

yes you are right, sorry...

OpenStudy (anonymous):

He spent $ 200.

OpenStudy (anonymous):

no problem, so first expression is going to be "x-5"

OpenStudy (turingtest):

right so we need to represent 200 by two variables, the one for price and the one for number of chairs. can you do that? 200=???

OpenStudy (turingtest):

no not x-5 first because we need two variables here to represent the two different things.

OpenStudy (anonymous):

x for chairs and y for 200

OpenStudy (turingtest):

Not quite, we will represent 200=Px where P is the price per chair and x is the number of chairs. Make sense?

OpenStudy (anonymous):

ok so, x-5 = Px?

OpenStudy (turingtest):

not quite... remember he sold the number of chars after 5 were stolen for 4 MORE than the original price P. so what does Px become? (both P and x change a bit)

OpenStudy (anonymous):

x-5 = Px +4 But then where did 200 go?

OpenStudy (turingtest):

close... 200=Px selling them we are told he changed the price by +4 and the thief changed the number of chairs by -5. We also know he MADE $40 total. So we have purchase: 200=Px selling: 200+40=(P+4)(x-5) reasonable? can you solve those two equations?

OpenStudy (turingtest):

remember (price per chair)(# of chairs)=(money spent/made)

OpenStudy (anonymous):

I cannot solve th eequation actually, i 'm stuck on 4x+px-5p-260=0

OpenStudy (turingtest):

Let me try...

OpenStudy (anonymous):

The answer at the end of the book is 25 and 8.

OpenStudy (anonymous):

he bought 25 chairs And the Cost of each chair was $8

OpenStudy (turingtest):

yes that is the solution that makes sense, now to prove it to you... it'll take a little work

OpenStudy (turingtest):

sales: 200+40=(p+4)(x-5) remember that 200=Px: Px+40=Px+4x-5P-20 60=4x-5P and we can write P=200/x subbing this in gives 60=4x-5(200/x)=4x-1000/x 60x=4x^2-1000 0=4x^2-60x-1000 can you solve it now? you will get two solutions, one will be negative and therefor make no sense. Throw that one away and keep the non-negative one.

OpenStudy (turingtest):

I think I made an error...

OpenStudy (turingtest):

yes, sorry, what did I do? the equations are correct originally I checked...

OpenStudy (anonymous):

Yes, I can!! Thank you so much TT for helping and explaining each and everything so elaborately, and for taking out time for me. I hope i can give you a medal for each of your reply. and each of the secon you spend at the post.

OpenStudy (turingtest):

Ok I wish I knew what I did wrong If yo put the equations 240=(p+4)(x-5) and 200=Px into wolfram you can see you get the right answer. http://www.wolframalpha.com/input/?i=200%3Dpx%26%26240%3D%28p%2B4%29%28x-5%29 So I guess good luck figuring out my mistake As for the last problem I don't think I can solve it... Hard!

OpenStudy (wasiqss):

the fraction is 3/7